OpenStudy (destinym823):

Find the equation of the circle that has a diameter with endpoints located at (7, 3) and (7, –5).

OpenStudy (evoker):

You need to find the center of the circle using the midpoint formula, and find the radius using the distance formula, and then finally (x-h)^2+(y-k)^2=r^2.

OpenStudy (sshayer):

eq. of circle with end points of diameter is (x-x1)(x-x2)+(y-y1)(y-y2)=0

OpenStudy (sshayer):



OpenStudy (destinym823):

would it be (x+1) squared +(y-7) squared =16 ?

OpenStudy (evoker):

close, your 1 and -7 are reversed

OpenStudy (destinym823):

oh my bad

OpenStudy (destinym823):

thanks

OpenStudy (sshayer):

\[(x-7)(x-7)+(y-3)(y+5)=0\] simplify it ,you will get the same eq.